Professor Wendy Smith has been offered the following deal: A law firm would like to retain her for an upfront payment of $ 60,000. In return, for the next year, the firm would have access to eight hours of her time every month. Smith's rate is $ 630 per hour, and her opportunity cost of capital is 14 % (equivalent annual rate, EAR).

What is the IRR (annual)? What does the IRR rule advice regarding this opportunity?

What is the NPV? What does the NPV rule say about this opportunity?

What is the IRR (annual)? The IRR (annual) is ________________%. What does the IRR rule advice regarding this opportunity?

Smith's cost of capital is 14 %, so according to the IRR rule, she should turn down this opportunity. What is the NPV?

The NPV is $_______________ What does the NPV rule say about this opportunity? The NPV is _______________,
Answer #1

Calculating NPV:

Present value at Time 0 = cashflow/(1+rate)time  

ra = 14% annually. We should convert it to monthly interest rate.

rm =14%/12 =1.17%

cash outflow of each month = -630*8 = -5040

Time(months) Cash flow PV of Cash flows
0 60000 60000
1 -5040 -4982
2 -5040 -4924
3 -5040 -4868
4 -5040 -4812
5 -5040 -4756
6 -5040 -4701
7 -5040 -4647
8 -5040 -4593
9 -5040 -4540
10 -5040 -4488
11 -5040 -4436
12 -5040 -4385
Sum 3867

Given, the sum of present value of cash flows is positive, the professor should accept the offer.

IRR calculation:

The rate at which the NPV becomes 0.

60000 = (5040/(1 + 1RR)! ) + (5040/(1-IRR)2 ) + (5040/ (1+ IRM ) + (5040/(1IRR))(5040/(1IRR)) (5040 IRR))(5040/1+ IRR))(5040

By hit and trial method, at IRR=0.12% monthly, the NPV is 0. As IRR< rm; the professor should reject the offer.